Alfagg:

The double number line shows the approximate number of ounces in a certain number of pounds: Based on the number line, about how many ounces are there in 8 pounds? 128 ounces 144 ounces 160 ounces 176 ounces

Alfagg:

here is the number line

1 attachment
Sock:

Ok so we to extend the number line, do you know how to do this?

Alfagg:

um.....

Sock:

ok I'll explain it to your then

Sock:

what is the difference between each number on the number in the ounces line

Sock:

ok so for ounces each plot goes up 16 everytime

Sock:

so we would expand the plot by adding 16 to the last number they gave us

Sock:

88 + 16 = ?

Alfagg:

104!

Sock:

ok good

Sock:

now for pounds to go up that would be easy because there only 1 apart, so what 6 + 1

Alfagg:

7

Sock:

good now we've got to go up each number line one more time

Sock:

so whats 104 + 16

Alfagg:

120

Sock:

good and 7+1

Alfagg:

8

Sock:

good now lets look at our numberline now

Sock:

1 attachment
Sock:

and as u can tell 8 pounds lands in the middle of 120 and the next number on the line

Sock:

so we have to divide 16 then add that total to 120, then we will have our answer

Sock:

so whats 16 divided by 2

Sock:

to write it out in an equation you would do 16 / 2 = x x + 120 = ?

Sock:

u there @Alfagg

Alfagg:

yeah

Alfagg:

i am in a zoom

Sock:

oh ok

Sock:

so u know what 16 / 2 is or?

Sock:

you know what its a simple question, I'll just give you the answer 16 / 2 = 8

Sock:

then you just have to add 120 + 8 and then you have your answer

Sock:

and 120 + 8 = 128

Sock:

so that is your answer :)

Sock:

hope I helped, buh-bye :D